OpenStudy (anonymous):

How many years will it take for a $50 investment to grow to $100 if it is compounded continuously at a rate of 5%? Given formula is: A(t)=P*e^rt Where P is initial investment, r is rate, t is time.

OpenStudy (anonymous):

100 = 50*e^0.05t 100/50 = e^0.05t 2= e^0.05t ln 2 = 0.05t t = (ln2)/0.05 = 13.86 years nearly. so 14 years i guess

OpenStudy (cwrw238):

A(t) = 50* e^(0.05*t) - like rkothaka said

OpenStudy (anonymous):

btw the $50 and the $100 have nothing to do with it. this is the "doubling time" so no matter what you start with it will double in \[\frac{\ln(2)}{.05}\] years